# Rectification from band gap oscillation

### v1

@@ I Introduction
The periodic metal-insulator transitions can potentially be used for voltage rectification purposes. We know that electrical transformers [20] use rectification properties of $p-n$ junction diodes [20] for converting alternating currents to direct currents which is required for running our electrical appliances. A $p-n$ junction is made from a $p$-type semiconductor which is basically hole doped and an $n$-type semiconductor which is electron doped. At their junction a depletion region is formed which decreases in width when the device is forward biased as opposed to the increasing width when the device is reverse biased. Ideally, therefore the device has very low resistance when forward biased and very high resistance when reverse biased and hence a single diode acts as a half-wave rectifier to an alternating voltage. We propose that a system periodically stretched and compressed with frequency $\omega$ can serve as a half-wave rectifier for an alternating voltage with same frequency $\omega$. In the metallic phase cycle, the system conducts while in the band insulator cycle the system does not conduct any electricity. Therefore, $p-n$ junction diodes in transformers can be replaced by such systems which can harbor a periodic metal-insulator transition in response to periodic pressure.
@@ III Rectification
Now if we consider an alternating electric field or equivalently a voltage which has the same time period as the periodic pressure on the chain then the half-filled chain would conduct in the half-cycle where it is a metal and will not conduct in the other half-cycle. This is shown in Fig 3$(a),(b)$. So, just like a p-n junction diode it acts like a half-wave rectifier, only the principles of operation are very different.
@@ IV Conclusion
In conclusion, we have proposed a one dimensional model of atoms in a chain whose length is being stretched and compressed periodically for example by providing oscillating pressure. As a consequence we see, that it is possible for the system to undergo repeated metal insulator transitions which can be used for the purpose of half-wave rectification of alternating voltages. Therefore it can potentially replace $p-n$ junction diodes whose rectification properties are used in transformers to convert alternating current to direct current. Moreover, our theory can be extended to systems where electron-electron interaction effects can not be neglected, for example in Mott insulators. Stretching reduces hopping and the electron -electron interaction and hopping ratio, $U/t$ becomes large and the system behaves like a Mott insulator at half-filling. While if $t$ become large i.e., if we compress the system, it will behave like a metal at half-filling. So, in systems where quasiparticle picture holds good and band theory works as well as in systems where electron-electron interaction plays a crucial role, periodic stretching and compression can be potentially used for rectification purposes, as in transformers replacing the widespread use of $p-n$ junction diodes.

### v2

@@ I Introduction
The periodic metal-insulator transitions can potentially be used for voltage rectification purposes. We know that electrical transformer [20] circuits use the rectification properties of $p-n$ junction diodes [20] for converting alternating currents to direct currents which is required for running our electrical appliances. A $p-n$ junction is made from a $p$-type semiconductor which is basically hole doped and an $n$-type semiconductor which is electron doped. At their junction a depletion region is formed which decreases in width when the device is forward biased as opposed to the increasing width when the device is reverse biased. Ideally, therefore the device has very low resistance when forward biased and very high resistance when reverse biased and hence a single diode acts as a half-wave rectifier to an alternating voltage. We propose that a system periodically stretched and compressed with frequency $\omega$ can serve as a half-wave rectifier for an alternating voltage with same frequency $\omega$. In the metallic phase cycle, the system conducts while in the band insulator cycle the system does not conduct any electricity. Therefore, $p-n$ junction diodes in transformer circuits can be replaced by such systems which can harbor a periodic metal-insulator transition in response to periodic pressure. One of the potential candidate materials is $TiS_{2}$ which shows a semiconductor to semimetal transition under pressure [21] that makes it worth exploring it’s behavior under time dependent periodic pressure.
@@ III Rectification
Now if we consider an alternating electric field or equivalently a voltage which has the same time period as the periodic pressure on the chain then the half-filled chain would conduct in the half-cycle where it is a metal and will not conduct in the other half-cycle. This is shown in Fig 3$(a),(b)$. So, just like a $p-n$ junction diode it acts like a half-wave rectifier, only the principles of operation are very different. However, at finite but low temperatures of operation, an asymmetry in the metallic and insulator regions is expected because for very small positive band gaps, thermal excitations can make the system metallic.
@@ IV Conclusion
In conclusion, we have proposed a one dimensional model of atoms in a chain whose length is being stretched and compressed periodically for example by providing oscillating pressure. As a consequence we see, that it is possible for the system to undergo repeated metal insulator transitions which can be used for the purpose of half-wave rectification of alternating voltages. Therefore it can potentially replace $p-n$ junction diodes whose rectification properties are used in transformer circuits to convert alternating current to direct current. Moreover, our theory can be extended to systems where electron-electron interaction effects can not be neglected, for example in Mott insulators. Stretching reduces hopping and the electron -electron interaction and hopping ratio, $U/t$ becomes large and the system behaves like a Mott insulator at half-filling. While if $t$ become large i.e., if we compress the system, it will behave like a metal at half-filling. So, in systems where quasiparticle picture holds good and band theory works as well as in systems where electron-electron interaction plays a crucial role, periodic stretching and compression can be potentially used for rectification purposes, as in transformer circuits, replacing the widespread use of $p-n$ junction diodes.